# Spectral selectors on lens spaces and applications to the geometry of the group of contactomorphisms

### v1

@@ 1. Introduction
The map $\nu:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\frac{2\pi}{k}\cdot\mathbb{Z}$ defined by
$\nu(\widetilde{\phi})=\max\left\{\,\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}\,,\,-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}\,\right\}$
is a stably unbounded conjugation invariant pseudonorm, which is compatible with the partial order $\leq$ and satisfies $\nu(\widetilde{\phi})\leq\frac{2\pi}{k}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})$ for every $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. The induced pseudonorm $\nu_{\ast}$ on $\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$ is non-degenerate and bounded.
Finally, we remark that the spectral selectors of Theorem 1.1 can also be used as in [4] to define a time function on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$, i.e. a function
$\tau:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\mathbb{R}$
@@
where $(\widetilde{\psi}_{j})_{j\geq 1}$ is any dense sequence in $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ with respect to the $\mathcal{C}^{1}$-topology. As in [4], the time function $\tau$ satisfies moreover $\tau(\widetilde{r_{T}}\cdot\widetilde{\phi})=T+\tau(\widetilde{\phi})$ for all $T$ and $\widetilde{\phi}$.
@@ Acknowledgments
The third author thanks Alberto Abbondandolo for the invitation to Bochum and kind hospitality in the fall of 2022, when part of this work has been done. We all thank Alberto also for his support and for inspiring discussions.
@@ 2. The non-linear Maslov index
As in [16], we first define the non-linear Maslov index assuming that $k$ is prime and then obtain the general case (Section 2) by pullback. Assume thus for now that $k$ is prime.
@@
Any contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity can be uniquely lifted to a $\mathbb{Z}_{k}$-equivariant contact isotopy $\{\bar{\phi}_{t}\}_{t\in[0,1]}$ of $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$ starting at the identity, which in turn can be uniquely extended to a conical Hamiltonian isotopy $\{\Phi_{t}\}_{t\in[0,1]}$ of $(\mathbb{R}^{2n},\omega_{0})$, i.e. a 1-parameter family of $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant homeomorphisms of $\mathbb{R}^{2n}$ that is a Hamiltonian isotopy on $\mathbb{R}^{2n}\smallsetminus\{0\}$. If $M$ is a multiple of $n$ then we say that a function $F:\mathbb{R}^{2M}\rightarrow\mathbb{R}$ is conical if it is $\mathcal{C}^{1}$ with Lipschitz differential, homogeneous of degree 2 with respect to the radial action of $\mathbb{R}_{>0}$ on $\mathbb{R}^{2M}$, and invariant by the diagonal action of $\mathbb{Z}_{k}$ on $\mathbb{R}^{2M}$. We say that a conical function $F:E\rightarrow\mathbb{R}$ defined on the total space of a trivial vector bundle $E=\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}^{2n}$ is a conical generating functions of a conical symplectomorphism $\Phi$ of $(\mathbb{R}^{2n},\omega_{0})$, i.e. a $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant homeomorphism of $\mathbb{R}^{2n}$ that is a symplectomorphism on $\mathbb{R}^{2n}\smallsetminus\{0\}$, if it is smooth near its fibre critical points other than the origin, $dF:E\rightarrow T^{\ast}E$ is transverse to the fibre conormal bundle $N_{E}^{\ast}$ except possibly at the origin, and $i_{F}$ is a homeomorphism between $\Sigma_{F}$ and the image of the graph of $\Phi$ by $\tau$. We say that $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$, $t\in[0,1]$, is a family of conical generating functions for a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity if for every $t$ the function $F_{t}$ is a conical generating function of $\Phi_{t}$, where $\{\Phi_{t}\}$ denotes the conical Hamiltonian isotopy of $(\mathbb{R}^{2n},\omega_{0})$ lifting $\{\phi_{t}\}$, and the map $(e,t)\mapsto F_{t}(e)$ is $\mathcal{C}^{1}$ with locally Lipschitz differential and smooth near $(e,t)$ whenever $e$ is a fibre critical point of $F_{t}$ other than the origin. We say that $F_{t}$, $t\in[0,1]$, is a based family of conical generating functions for $\{\phi_{t}\}_{t\in[0,1]}$ if moreover $F_{0}$ is equivalent to the zero function on $\mathbb{R}^{2n}$, where we consider on the set of conical generating functions the smallest equivalence relation under which two such functions are equivalent if they differ by a stabilization (i.e. replacing a conical generating function $F:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ by $F\oplus Q:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\times\mathbb{R}^{2nN^{\prime}}\rightarrow\mathbb{R}$ for a non-degenerate $\mathbb{Z}_{k}$-invariant quadratic form $Q$ on $\mathbb{R}^{2nN^{\prime}}$) or by a fibre preserving conical homeomorphism (i.e. a $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant homeomorphism of $\mathbb{R}^{2n}\times\mathbb{R}^{2nN}$ that takes each fibre $\{z\}\times\mathbb{R}^{2nN}$ to itself) that restricts to a diffeomorphism between neighborhoods of fibre critical points other than the origin. It is proved in [16, Proposition 2.14] that any contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity has a based family $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ of conical generating functions.
@@
The following property is proved in [16, Corollary 3.15] (cf [16, Proposition 3.14, Proposition 3.9 (v) and Remark 3.11] for the case $k=2$).
@@
where $F_{t}$, $t\in[0,1]$, is any based family of conical generating functions for $\{\phi_{t}\}_{t\in[0,1]}$. It is proved in [16, Proposition 2.20] that any two based families of conical generating functions for $\{\phi_{t}\}_{t\in[0,1]}$ are equivalent, where we consider on the set of based families of conical generating functions the smallest equivalence relation under which two such families are equivalent if they differ by a stabilization (i.e. replacing $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ by $F_{t}\oplus Q:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\times\mathbb{R}^{2nN^{\prime}}\rightarrow\mathbb{R}$ for a non-degenerate $\mathbb{Z}_{k}$-invariant quadratic form $Q$ on $\mathbb{R}^{2nN^{\prime}}$) or by a 1-parameter family of fibre preserving conical homeomorphism that restrict to diffeomorphisms between neighborhoods of fibre critical points other than the origin. Since, for $k>2$, $\operatorname{ind}(Q)$ is even for every $\mathbb{Z}_{k}$-invariant quadratic form $Q$ ([16, Remark 3.13]), it thus follows from Section 2 that $\mu(\{\phi_{t}\}_{t\in[0,1]})$ is well-defined, i.e. it does not depend on the choice of a based family of conical generating functions. Moreover, it is proved in [16] (as a consequence of [16, Proposition 2.21]) that $\mu$ descends to a map
@@
By definition we have $\mu(\widetilde{\mathrm{id}})=0$. By [16, Example 4.1], if $\widetilde{\phi}$ is small enough in the $\mathcal{C}^{1}$-topology then $0\leq\mu(\widetilde{\phi})\leq 2n$, and if moreover $\widetilde{\phi}$ is positive then $\mu(\widetilde{\phi})=2n$. In particular, for $\epsilon>0$ small enough we have $\mu(\widetilde{r_{\epsilon}})=2n$. By [16, Example 4.13], for every integer $m$ we have $\mu(\widetilde{r_{2\pi m}})=2nm$.
It is proved in [16, Theorem 1.4 (i) and Remark 1.7] that for any two elements $\widetilde{\phi}$ and $\widetilde{\psi}$ of $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ we have
@@
Before proving Section 2, recall ([16, Proposition 2.10]) that if $F\colon\mathbb{R}^{2n}\times\mathbb{R}^{2nN_{1}}\to\mathbb{R}$ and $G\colon\mathbb{R}^{2n}\times\mathbb{R}^{2nN_{2}}\to\mathbb{R}$ are conical generating functions for conical symplectomorphisms $\Phi$ and $\Psi$ respectively, then the function $F\,\sharp\,G\colon\mathbb{R}^{2n}\times(\mathbb{R}^{2n}\times\mathbb{R}^{2n}\times\mathbb{R}^{2nN_{1}}\times\mathbb{R}^{2nN_{2}})\rightarrow\mathbb{R}$ defined by
@@
is a conical generating function of $\Psi\circ\Phi$, and ([16, Proposition 2.26]) there is a linear $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant injection
@@
such that $(F\,\sharp\,G)\circ\,\iota=F\oplus G$. Since the cohomological index is monotone, i.e. $\operatorname{ind}(A)\leq\operatorname{ind}(B)$ if $A\subset B$ [16, Proposition 3.9 (i)], we deduce that
@@
Recall also ([16, Lemma 4.2]) that if $F$ and $G$ are equivalent to the zero function then
@@
Let $F_{t}$ and $G_{t}$ be based families of conical generating functions for contact isotopies $\{\phi_{t}\}_{t\in[0,1]}$ and $\{\psi_{t}\}_{t\in[0,1]}$ representing $\widetilde{\phi}$ and $\widetilde{\psi}$ respectively. Then, by [16, Proposition 2.10 and Remark 2.14], $G_{t}\,\sharp\,F_{t}$ is a based family of conical generating functions for $\{\phi_{t}\circ\psi_{t}\}_{t\in[0,1]}$, and thus
@@
if either $k=2$ or $\operatorname{ind}(F_{1})$ is even or $\operatorname{ind}(G_{1})$ is even, hence (since, for $k>2$, $\operatorname{ind}(F_{0})$ and $\operatorname{ind}(G_{0})$ are even by [16, Remark 3.13]) if either $k=2$ or $\mu(\widetilde{\phi})$ is even or $\mu(\widetilde{\psi})$ is even. ∎
@@
if either $\widetilde{\phi}$ or $\widetilde{\psi}$ are in $\pi_{1}\big{(}\operatorname{Cont}_{0}(\mathbb{RP}^{2n-1},\xi_{0})\big{)}$. For our applications we only need this property in the case when one of the factors is the Reeb flow.
@@
where $\{\phi_{t}\}_{t\in[0,1]}$ is a contact isotopy representing $\widetilde{\phi}$, and we consider a based family $F_{t}$, $t\in[0,1]$, of conical generating functions for $\{\varphi_{t}\}_{t\in[0,1]}$ so that $F_{t}$, $t\in[0,\frac{1}{2}]$, is a family of quadratic forms generating $\{r_{4\pi mt}\}_{t\in[0,\frac{1}{2}]}$ (cf. [16, Proposition 4.9]). Since $\{r_{4\pi mt}\}_{t\in[0,\frac{1}{2}]}$ is a loop, by [16, Lemma 4.10] the quadratic form $F_{\frac{1}{2}}$ is equivalent to the zero function, and so
@@
Before proving Section 2, recall that the cohomological index is continuous ([16, Proposition 3.9 (ii)]): every closed subset $A$ of $L^{2M-1}_{k}$ has a neighborhood $\mathcal{U}$ such that if $A\subset\mathcal{V}\subset\mathcal{U}$ then $\operatorname{ind}(\mathcal{V})=\operatorname{ind}(A)$.
@@
By monotonicity of generating functions ([16, Proposition 2.23]) we can assume that $\frac{df_{t}}{dt}\leq 0$ for all $t\in[\frac{1}{2},1]$, and so
@@
It is proved in [16, Proposition 4.21] that if $\{\phi_{t}\}_{t\in[0,1)}$ is a non-negative (respectively non-positive) contact isotopy then $\mu([\{\phi_{t}\}_{t\in[0,1)}])\geq 0$ (respectively $\mu([\{\phi_{t}\}_{t\in[0,1)}])\leq 0$). We actually have the following result.
@@
of a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ representing $\widetilde{\phi}$ and a non-negative contact isotopy $\{\chi_{t}\}_{t\in[0,1]}$. Let $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ be a based family of generating functions for $\{\psi_{t}\}_{t\in[0,1]}$. By monotonicity of generating function ([16, Proposition 2.23]) we can assume that $F_{1}\geq F_{\frac{1}{2}}$and so, by monotonicity of the cohomological index ([16, Proposition 3.9 (i)]), $\operatorname{ind}(F_{1})\leq\operatorname{ind}(F_{\frac{1}{2}})$. We thus have
@@
If $F_{t}$ is a based family of conical generating functions for a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ then for every $t$ there is a 1–1 correspondence between the critical points of critical value zero of $f_{t}$ and the discriminant points of $\phi_{t}$ that lift to discriminant points of $\bar{\phi}_{t}$, where $\{\bar{\phi}_{t}\}_{t\in[0,1]}$ is the lift of $\{\phi_{t}\}_{t\in[0,1]}$ to $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$. Since the non-linear Maslov index $\mu(\{\phi_{t}\}_{t\in[0,1]})$ counts, with multiplicity given by the change in the cohomological index of the sublevel sets $\{f_{t}\leq 0\}$, the critical points of $f_{t}$ with critical value zero as $t$ varies in $[0,1]$, its value is related to the presence of discriminant points of $\bar{\phi}_{t}$ for $t\in[0,1]$. More precisely, we have the following result ([16, Theorem 1.4 (iii)]).
@@
Let $\{\phi_{t}\}_{t\in[0,1]}$ be a contact isotopy of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity, and let $[t_{0},t_{1}]$ be a subinterval of $[0,1]$. If $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})\neq\mu(\{\phi_{t}\}_{t\in[0,t_{1}]})$ then there is $\underline{t}\in[t_{0},t_{1}]$ such that $\bar{\phi}_{\underline{t}}$ has discriminant points. If moreover [^1] In [16, Theorem 1.4 (iii)] it is assumed that there is only one $\underline{t}\in[0,1]$ such that $\bar{\phi}_{\underline{t}}$ has discriminant points. However, the proof works in the same way also if we only assume that $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $[t_{0},\underline{t})$ and on $(\underline{t},t_{1}]$. Similarly for Section 2 below. the map $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $[t_{0},\underline{t})$ and on $(\underline{t},t_{1}]$ and all the discriminant points of $\bar{\phi}_{\underline{t}}$ are non-degenerate then

### v2

@@ 1. Introduction
The map $\nu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow T_{\underline{w}}\cdot\mathbb{Z}$ defined by
$\nu(\widetilde{\phi})=\max\left\{\,\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}\,,\,-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}\,\right\}$
is a stably unbounded conjugation invariant pseudonorm, which is compatible with the partial order $\leq$ and satisfies $\nu(\widetilde{\phi})\leq T_{\underline{w}}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})$ for every $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. The induced pseudonorm $\nu_{\ast}$ on $\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ is non-degenerate and bounded.
Finally, we remark that the spectral selectors of Theorem 1.1 can also be used as in [4] to define a time function on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$, i.e. a function
$\tau:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathbb{R}$
@@
where $(\widetilde{\psi}_{j})_{j\geq 1}$ is any dense sequence in $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ with respect to the $\mathcal{C}^{1}$-topology. As in [4], the time function $\tau$ satisfies moreover $\tau(\widetilde{r_{T}}\cdot\widetilde{\phi})=T+\tau(\widetilde{\phi})$ for all $T$ and $\widetilde{\phi}$.
@@ Acknowledgments
The third author thanks Alberto Abbondandolo for the invitation to Bochum and kind hospitality in the fall of 2022, when part of this work has been done. We all thank Alberto also for his support and for inspiring discussions. The version of this article that was first posted to arXiv contained a mistake concerning the period of the Reeb flow on lens spaces with general weights and affecting in particular Section 1, which was stated in more generality than what we can actually prove. We are grateful to Miguel Abreu for noticing this mistake and for very useful discussions.
@@ 2. The non-linear Maslov index
$\mu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathbb{Z}$
@@
Since the weights $\underline{w}$ do not play a particular role in the discussion, in this section we denote $L_{k}^{2n-1}(\underline{w})$ simply by $L_{k}^{2n-1}$. As in [17], we first define the non-linear Maslov index assuming that $k$ is prime and then obtain the general case (Section 2) by pullback. Assume thus for now that $k$ is prime.
@@
Any contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity can be uniquely lifted to a $\mathbb{Z}_{k}$-equivariant contact isotopy $\{\bar{\phi}_{t}\}_{t\in[0,1]}$ of $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$ starting at the identity, which in turn can be uniquely extended to a conical Hamiltonian isotopy $\{\Phi_{t}\}_{t\in[0,1]}$ of $(\mathbb{R}^{2n},\omega_{0})$, i.e. a 1-parameter family of $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant homeomorphisms of $\mathbb{R}^{2n}$ that is a Hamiltonian isotopy on $\mathbb{R}^{2n}\smallsetminus\{0\}$. If $M$ is a multiple of $n$ then we say that a function $F:\mathbb{R}^{2M}\rightarrow\mathbb{R}$ is conical if it is $\mathcal{C}^{1}$ with Lipschitz differential, homogeneous of degree 2 with respect to the radial action of $\mathbb{R}_{>0}$ on $\mathbb{R}^{2M}$, and invariant by the diagonal action of $\mathbb{Z}_{k}$ on $\mathbb{R}^{2M}$. We say that a conical function $F:E\rightarrow\mathbb{R}$ defined on the total space of a trivial vector bundle $E=\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}^{2n}$ is a conical generating functions of a conical symplectomorphism $\Phi$ of $(\mathbb{R}^{2n},\omega_{0})$, i.e. a $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant homeomorphism of $\mathbb{R}^{2n}$ that is a symplectomorphism on $\mathbb{R}^{2n}\smallsetminus\{0\}$, if it is smooth near its fibre critical points other than the origin, $dF:E\rightarrow T^{\ast}E$ is transverse to the fibre conormal bundle $N_{E}^{\ast}$ except possibly at the origin, and $i_{F}$ is a homeomorphism between $\Sigma_{F}$ and the image of the graph of $\Phi$ by $\tau$. We say that $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$, $t\in[0,1]$, is a family of conical generating functions for a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity if for every $t$ the function $F_{t}$ is a conical generating function of $\Phi_{t}$, where $\{\Phi_{t}\}$ denotes the conical Hamiltonian isotopy of $(\mathbb{R}^{2n},\omega_{0})$ lifting $\{\phi_{t}\}$, and the map $(e,t)\mapsto F_{t}(e)$ is $\mathcal{C}^{1}$ with locally Lipschitz differential and smooth near $(e,t)$ whenever $e$ is a fibre critical point of $F_{t}$ other than the origin. We say that $F_{t}$, $t\in[0,1]$, is a based family of conical generating functions for $\{\phi_{t}\}_{t\in[0,1]}$ if moreover $F_{0}$ is equivalent to the zero function on $\mathbb{R}^{2n}$, where we consider on the set of conical generating functions the smallest equivalence relation under which two such functions are equivalent if they differ by a stabilization (i.e. replacing a conical generating function $F:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ by $F\oplus Q:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\times\mathbb{R}^{2nN^{\prime}}\rightarrow\mathbb{R}$ for a non-degenerate $\mathbb{Z}_{k}$-invariant quadratic form $Q$ on $\mathbb{R}^{2nN^{\prime}}$) or by a fibre preserving conical homeomorphism (i.e. a $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant homeomorphism of $\mathbb{R}^{2n}\times\mathbb{R}^{2nN}$ that takes each fibre $\{z\}\times\mathbb{R}^{2nN}$ to itself) that restricts to a diffeomorphism between neighborhoods of fibre critical points other than the origin. It is proved in [17, Proposition 2.14] that any contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity has a based family $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ of conical generating functions.
@@
The following property is proved in [17, Corollary 3.15] (cf [17, Proposition 3.14, Proposition 3.9 (v) and Remark 3.11] for the case $k=2$).
@@
where $F_{t}$, $t\in[0,1]$, is any based family of conical generating functions for $\{\phi_{t}\}_{t\in[0,1]}$. It is proved in [17, Proposition 2.20] that any two based families of conical generating functions for $\{\phi_{t}\}_{t\in[0,1]}$ are equivalent, where we consider on the set of based families of conical generating functions the smallest equivalence relation under which two such families are equivalent if they differ by a stabilization (i.e. replacing $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ by $F_{t}\oplus Q:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\times\mathbb{R}^{2nN^{\prime}}\rightarrow\mathbb{R}$ for a non-degenerate $\mathbb{Z}_{k}$-invariant quadratic form $Q$ on $\mathbb{R}^{2nN^{\prime}}$) or by a 1-parameter family of fibre preserving conical homeomorphism that restrict to diffeomorphisms between neighborhoods of fibre critical points other than the origin. Since, for $k>2$, $\operatorname{ind}(Q)$ is even for every $\mathbb{Z}_{k}$-invariant quadratic form $Q$ ([17, Remark 3.13]), it thus follows from Section 2 that $\mu(\{\phi_{t}\}_{t\in[0,1]})$ is well-defined, i.e. it does not depend on the choice of a based family of conical generating functions. Moreover, it is proved in [17] (as a consequence of [17, Proposition 2.21]) that $\mu$ descends to a map
@@
By definition we have $\mu(\widetilde{\mathrm{id}})=0$. By [17, Example 4.1], if $\widetilde{\phi}$ is small enough in the $\mathcal{C}^{1}$-topology then $0\leq\mu(\widetilde{\phi})\leq 2n$, and if moreover $\widetilde{\phi}$ is positive then $\mu(\widetilde{\phi})=2n$. In particular, for $\epsilon>0$ small enough we have $\mu(\widetilde{r_{\epsilon}})=2n$. By [17, Example 4.13], for every integer $m$ we have $\mu(\widetilde{r_{2\pi m}})=2nm$.
It is proved in [17, Theorem 1.4 (i) and Remark 1.7] that for any two elements $\widetilde{\phi}$ and $\widetilde{\psi}$ of $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ we have
@@
Before proving Section 2, recall ([17, Proposition 2.10]) that if $F\colon\mathbb{R}^{2n}\times\mathbb{R}^{2nN_{1}}\to\mathbb{R}$ and $G\colon\mathbb{R}^{2n}\times\mathbb{R}^{2nN_{2}}\to\mathbb{R}$ are conical generating functions for conical symplectomorphisms $\Phi$ and $\Psi$ respectively, then the function $F\,\sharp\,G\colon\mathbb{R}^{2n}\times(\mathbb{R}^{2n}\times\mathbb{R}^{2n}\times\mathbb{R}^{2nN_{1}}\times\mathbb{R}^{2nN_{2}})\rightarrow\mathbb{R}$ defined by
@@
is a conical generating function of $\Psi\circ\Phi$, and ([17, Proposition 2.26]) there is a linear $(\mathbb{Z}_{k}\times\mathbb{R}_{>0})$-equivariant injection
@@
such that $(F\,\sharp\,G)\circ\,\iota=F\oplus G$. Since the cohomological index is monotone, i.e. $\operatorname{ind}(A)\leq\operatorname{ind}(B)$ if $A\subset B$ [17, Proposition 3.9 (i)], we deduce that
@@
Recall also ([17, Lemma 4.2]) that if $F$ and $G$ are equivalent to the zero function then
@@
Let $F_{t}$ and $G_{t}$ be based families of conical generating functions for contact isotopies $\{\phi_{t}\}_{t\in[0,1]}$ and $\{\psi_{t}\}_{t\in[0,1]}$ representing $\widetilde{\phi}$ and $\widetilde{\psi}$ respectively. Then, by [17, Proposition 2.10 and Remark 2.14], $G_{t}\,\sharp\,F_{t}$ is a based family of conical generating functions for $\{\phi_{t}\circ\psi_{t}\}_{t\in[0,1]}$, and thus
@@
if either $k=2$ or $\operatorname{ind}(F_{1})$ is even or $\operatorname{ind}(G_{1})$ is even, hence (since, for $k>2$, $\operatorname{ind}(F_{0})$ and $\operatorname{ind}(G_{0})$ are even by [17, Remark 3.13]) if either $k=2$ or $\mu(\widetilde{\phi})$ is even or $\mu(\widetilde{\psi})$ is even. ∎
@@
if either $\mathcal{L}(\widetilde{\phi})$ or $\mathcal{L}(\widetilde{\psi})$ are in $\pi_{1}\big{(}\operatorname{Cont}_{0}(\mathbb{S}^{2n-1},\bar{\xi}_{0})\big{)}$. For our applications we only need this property in the case when one of the factors is the Reeb flow.
@@
where $\{\phi_{t}\}_{t\in[0,1]}$ is a contact isotopy representing $\widetilde{\phi}$, and we consider a based family $F_{t}$, $t\in[0,1]$, of conical generating functions for $\{\varphi_{t}\}_{t\in[0,1]}$ so that $F_{t}$, $t\in[0,\frac{1}{2}]$, is a family of quadratic forms generating $\{r_{4\pi mt}\}_{t\in[0,\frac{1}{2}]}$ (cf. [17, Proposition 4.9]). Since the lift of $\{r_{4\pi mt}\}_{t\in[0,\frac{1}{2}]}$ to $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$ is a loop, by [17, Lemma 4.10] the quadratic form $F_{\frac{1}{2}}$ is equivalent to the zero function, and so
@@
Before proving Section 2, recall that the cohomological index is continuous ([17, Proposition 3.9 (ii)]): every closed subset $A$ of $L^{2M-1}_{k}$ has a neighborhood $\mathcal{U}$ such that if $A\subset\mathcal{V}\subset\mathcal{U}$ then $\operatorname{ind}(\mathcal{V})=\operatorname{ind}(A)$.
@@
By monotonicity of generating functions ([17, Proposition 2.23]) we can assume that $\frac{df_{t}}{dt}\leq 0$ for all $t\in[\frac{1}{2},1]$, and so
@@
It is proved in [17, Proposition 4.21] that if $\{\phi_{t}\}_{t\in[0,1)}$ is a non-negative (respectively non-positive) contact isotopy then $\mu([\{\phi_{t}\}_{t\in[0,1)}])\geq 0$ (respectively $\mu([\{\phi_{t}\}_{t\in[0,1)}])\leq 0$). We actually have the following result.
@@
of a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ representing $\widetilde{\phi}$ and a non-negative contact isotopy $\{\chi_{t}\}_{t\in[0,1]}$. Let $F_{t}:\mathbb{R}^{2n}\times\mathbb{R}^{2nN}\rightarrow\mathbb{R}$ be a based family of generating functions for $\{\psi_{t}\}_{t\in[0,1]}$. By monotonicity of generating function ([17, Proposition 2.23]) we can assume that $F_{1}\geq F_{\frac{1}{2}}$and so, by monotonicity of the cohomological index ([17, Proposition 3.9 (i)]), $\operatorname{ind}(F_{1})\leq\operatorname{ind}(F_{\frac{1}{2}})$. We thus have
@@
If $F_{t}$ is a based family of conical generating functions for a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ then for every $t$ there is a 1–1 correspondence between the critical points of critical value zero of $f_{t}$ and the discriminant points of $\phi_{t}$ that lift to discriminant points of $\bar{\phi}_{t}$, where $\{\bar{\phi}_{t}\}_{t\in[0,1]}$ is the lift of $\{\phi_{t}\}_{t\in[0,1]}$ to $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$. Since the non-linear Maslov index $\mu(\{\phi_{t}\}_{t\in[0,1]})$ counts, with multiplicity given by the change in the cohomological index of the sublevel sets $\{f_{t}\leq 0\}$, the critical points of $f_{t}$ with critical value zero as $t$ varies in $[0,1]$, its value is related to the presence of discriminant points of $\bar{\phi}_{t}$ for $t\in[0,1]$. More precisely, we have the following result ([17, Theorem 1.4 (iii)]).
@@
Let $\{\phi_{t}\}_{t\in[0,1]}$ be a contact isotopy of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity, and let $[t_{0},t_{1}]$ be a subinterval of $[0,1]$. If $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})\neq\mu(\{\phi_{t}\}_{t\in[0,t_{1}]})$ then there is $\underline{t}\in[t_{0},t_{1}]$ such that $\bar{\phi}_{\underline{t}}$ has discriminant points. If moreover [^1] In [17, Theorem 1.4 (iii)] it is assumed that there is only one $\underline{t}\in[0,1]$ such that $\bar{\phi}_{\underline{t}}$ has discriminant points. However, the proof works in the same way also if we only assume that $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $[t_{0},\underline{t})$ and on $(\underline{t},t_{1}]$. Similarly for Section 2 below. the map $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $[t_{0},\underline{t})$ and on $(\underline{t},t_{1}]$ and all the discriminant points of $\bar{\phi}_{\underline{t}}$ are non-degenerate then